# On the Importance of Uncertainty in Decision-Making with Large Language Models

### v1

@@ Abstract
We investigate the role of uncertainty in decision-making problems with natural language as input. For such tasks, using Large Language Models as agents has become the norm. However, none of the recent approaches employ any additional phase for estimating the uncertainty the agent has about the world during the decision-making task. We focus on a fundamental decision-making framework with natural language as input, which is the one of contextual bandits, where the context information consists of text. As a representative of the approaches with no uncertainty estimation, we consider an LLM bandit with a greedy policy, which picks the action corresponding to the largest predicted reward. We compare this baseline to LLM bandits that make active use of uncertainty estimation by integrating the uncertainty in a Thompson Sampling policy. We employ different techniques for uncertainty estimation, such as Laplace Approximation, Dropout, and Epinets. We empirically show on real-world data that the greedy policy performs worse than the Thompson Sampling policies. These findings suggest that, while overlooked in the LLM literature, uncertainty plays a fundamental role in bandit tasks with LLMs.
@@ 1 Introduction
In parallel, many critical real-world systems are increasingly relying on these models to make decisions (Yang et al., 2023b), where the consequences of a particular action are typically reflected by a reward signal. This usually works by fine-tuning an LLM to serve as a reward model, where we try to teach the network to predict the mean reward for each action. A common way of using this model to make decisions is to use the predictions greedily (Riquelme et al., 2018), taking the action with the greatest estimated reward. However, this approach ignores the fact that reward estimates may be inaccurate, which can lead to never-ending pathological behavior (Russo et al., 2018). On the other hand, the literature on contextual bandits (Wang et al., 2005) provides a principled approach to deal with inaccurate reward estimates in a decision-making task. A popular modelling tool for doing this in the Bayesian framework is to maintain a probabilistic reward model that relies on two separate notions of uncertainty: epistemic and aleatoric. Epistemic uncertainty is the one that reflects the fact that we have not yet seen enough fine-tuning data to estimate the mean reward well, while aleatoric uncertainty reflects the irreducible noise associated with observing a reward. When such a reward model is coupled with an action-selection algorithm like Thompson Sampling (Thompson, 1933), high-quality estimates of epistemic uncertainty can be leveraged to produce a better balance between exploration and exploitation in the decision-making task, evidenced by lower regret. In this paper, we want to investigate the importance of employing epistemic uncertainty models in decision-making tasks with natural language and LLMs. In particular, we focus on one of the most fundamental decision-making tasks with natural language as input: contextual bandits. In a contextual bandit problem, the agent must continually take actions based on observed context, which in our case is text, and observes the reward for the chosen action. This framework is applicable to a myriad of real-world scenarios. A concrete example application we will consider is the problem of automated content moderation on online platforms (Vaccaro et al., 2020; Gorwa et al., 2020; Ma & Kou, 2021). In such a scenario, we have users who may post on a platform, and an algorithm then decides whether or not the content is safe to publish. In this paper, we study this scenario empirically, combining the insights from the LLM and bandit communities. The greedy approach outlined above is our main baseline and was chosen for its simplicity and ease of implementation. We benchmark it against approximate Bayesian models coupled with the Thompson Sampling algorithm for action selection. Since Thompson Sampling requires epistemic uncertainty to be of high quality to work well, we investigate several different approaches of doing so. Specifically, we compare techniques for epistemic uncertainty estimation that can be scaled up to be used with current LLMs. These are the Laplace Approximation (Daxberger et al., 2021), Dropout (Gal & Ghahramani, 2016) and Epinets (Osband et al., 2023a). We show that all our TS policies significantly and consistently outperform the greedy baseline on a real-world bandit benchmark. A preview of the results is shown in Figure 1. These findings suggest that, while often overlooked, epistemic uncertainty plays a fundamental role in bandit tasks with LLMs.
@@
  • We provide a comprehensive empirical study of LLM-based contextual bandits;
@@
  • We empirically show on real-world data that the greedy policy is sub-optimal, and that epistemic uncertainty is fundamental for solving contextual bandit problems with text data.
@@ 2 Preliminaries
In this section, we describe the problem statement, and then we introduce the topic of uncertainty in Machine Learning models.
@@ 2.1 Batch Contextual Bandit problem
The contextual bandit problem is a sequential decision-making framework where an agent interacts with an environment over a number of time steps. In particular, we focus on the more general case where, for each time step, we observe a batch of contexts and the agent has to select a batch of actions (Kandasamy et al., 2018). Let us call the total number of time steps $T$. At each time step $t=1,2,\dots,T$, the agent observes a batch of contexts, which we denote with $x^{1}_{t},\dots,x^{B}_{t}$. Each context $x^{b}_{t}$ contains useful information that should be used to solve the task. In this paper, we focus on decision-making problems driven by text. Hence, in this case, each context $x^{b}_{t}$ consists of text information. The agent then uses the observed data until now, which we call $\mathcal{D}$, to select an action $a^{b}_{t}$ from a set ${\mathcal{A}}$ of $K$ possible actions, conditioned on the given context $x^{b}_{t}$ with $b\in\{1,\dots,B\}$. For each selected action $a^{b}_{t}$, the agent receives a reward $r^{b}_{t}$, that depends on the context $x^{b}_{t}$ and action $a^{b}_{t}$. At the end of the time step, we have observed a set of tuples that we denote as ${\mathcal{D}}_{t}=\{(x^{1}_{t},a^{1}_{t},r^{1}_{t}),\dots,(x^{B}_{t},a^{B}_{t},r^{B}_{t})\}$. This set of tuples is added to the observed dataset $\mathcal{D}\leftarrow\mathcal{D}\cup{\mathcal{D}}_{t}$. The aim of the agent is to learn a policy that minimizes the regret over the $T$ time steps. We define the average regret as the difference between the cumulative reward from the learned policy and the cumulative reward from an optimal policy, all divided by the number of time steps. We define the optimal policy $\pi^{*}$ as the policy that, within the considered policy space $\Pi$, maximizes the expected reward: $\pi^{*}=\arg\max_{\pi\in\Pi}\mathbb{E}_{x}\mathbb{E}_{a\sim\pi(\cdot\mid x)}\left[r(x,a)\right]$. Let us call $r^{*}_{t}$ the expected rewards obtained by the optimal policy at any time step $t$. We can define the average regret as follows:
@@ 3.1 Greedy policy
where $\lambda=\sigma^{2}_{\mathrm{obs}}/\sigma^{2}_{p}$, and we set $\theta^{(0)}=\theta_{p}$. This loss now only assumes that the data points in ${\mathcal{D}}_{t}$ are i.i.d., and we update the prior at each time step with the weights obtained at the previous time step. This will be the loss we will use to optimize the greedy bandit in our experimental analysis in Section 5.
@@ 3.2 Thompson Sampling
Thompson Sampling (TS) (Thompson, 1933; Russo et al., 2018) is a probabilistic algorithm for the contextual bandit problem. The key idea of Thompson Sampling is to maintain an epistemic uncertainty estimate in the form of a posterior distribution over the parameters of the model $P(\theta|{\mathcal{D}})$. Initially, this distribution is set to a prior distribution $P(\theta)$ that represents the agent’s initial uncertainty. On each time step, for each observed context, the agent samples a set of parameters $\hat{\theta}\sim P(\theta|{\mathcal{D}})$ from the posterior distribution, and selects the action with the highest reward according to the sampled model. At the end of the time step, the agent updates the posterior distribution of the parameters with the observed data. A high-level description of this procedure is provided in Algorithm 2. This approach balances exploration and exploitation by sampling from a posterior that quantifies the epistemic uncertainty. Actions with more uncertain estimates will be explored, while actions that are currently believed to have higher rewards will be exploited. TS can be thought as taking actions according to the (epistemic) probability that they are optimal, which leads to a good balance between exploration and exploitation. The posterior distribution concentrates over time, automatically adjusting the exploration-exploitation trade-off.
This approach has been shown to be effective both in theory and in practice (Chapelle & Li, 2011). When the Bayesian updates are exact, there are theoretical guarantees on the performance of TS agents (Agrawal & Goyal, 2017). However, in our case, we are dealing with a complex bandit model that uses a pre-trained LLM. Therefore, an exact Bayesian update is computationally infeasible, and we have to resort to approximations. In Section 4, we will show different techniques to estimate posterior distribution of the parameters of a neural network and how to adapt such techniques to our case of LLM bandits.
@@ 4.1 Dropout
The dropout technique (Srivastava et al., 2014) consists of randomly setting a proportion $p$ of neuron outputs to zero at each forward pass through the network during training. This random “dropping out” of neurons allows the network to sample and train on different (but overlapping) architectures. In standard supervised learning, dropout is then deactivated during inference, and all dropout neurons are re-scaled to account for the fact that all dropout neurons are active. However, in our case, we still apply dropout during the action selection (Gal & Ghahramani, 2016; Riquelme et al., 2018). Using dropout in this phase, we randomly select (according to the dropout probability) a set of parameters $\hat{\theta}$ to use. This procedure can be seen as sampling parameters from an approximate posterior distribution: $\hat{\theta}\sim P(\theta|{\mathcal{D}})$. With such approximate posterior distribution, we can apply Thompson Sampling. This technique is valuable for large models, such as our LLM bandit, because it does not add any overhead to the procedure and does not require additional memory.
@@ 4.2 Laplace Approximation
Notice that the first-order term vanishes because we are expanding around $\theta_{\mathrm{MAP}}$, which is a point of minimum of the loss. After some algebraic manipulations, it can be shown that the posterior distribution can be expressed as [^2] For a more detailed derivation, see Appendix A.:
@@
This approximation has the advantage of being positive semi-definite and (Kunstner et al., 2019) showed that it is accurate when the regression residuals are small. For more details on this approximation, see Appendix A.
@@ 4.3 Last-Layer Laplace Approximation
Alternatively, instead of using Hessian approximations, another way to scale LA for large models is to compute the full Hessian but only for a subset of the parameters of the network, which typically is the last layer of the model. In our case, this means that we will compute the Hessian for the randomly initialized final layer, while the pre-trained LLM layers remain fixed during the sampling phase of TS (every parameter is still fine-tuned during training). This approximation can reduce the exploration due to the fixed weights, but at the same time it allows a full quadratic Hessian computation, leading to a better quality of exploration. Last layer LA is also equivalent to Bayesian linear regression (Box & Tiao, 2011) on the features before the last layer within a single batch. Also, with Gaussian likelihood, last-layer LA is exact within a single batch.
@@ 5 Experiments
In this section, we provide a comprehensive empirical study of LLM-based contextual bandits on real-world data. We show how to adapt the epistemic uncertainty estimation techniques we described to bandits with pre-trained LLMs. In particular, we identify the best variant of the Laplace approximation and dropout to use, in addition to a proposed Epinet architecture that we show works well for TS with LLM bandits. Finally, we empirically show that the greedy policy is sub-optimal, shedding light on the importance of epistemic uncertainty for LLM bandits.
@@ 5.1 Experimental Methodology
Task
We evaluate the bandit policies on an automated content moderation task. We use an open-source dataset, called “measuring hate speech”, which is openly available on HuggingFace Datasets [^4] https://huggingface.co/datasets/ucberkeley-dlab/measuring-hate-speech. This dataset consists of around 136,000 comments. Each of them is associated with a continuous score, called the “hate speech score”, provided for each comment. Comments with a score $>0.5$ are considered “toxic” (around $36\%$ of the comments are labeled as toxic, while the others are not toxic). The task is framed as a contextual bandit problem where the context is the text comment, and the actions available to the agent are “not publish” or “publish”. For each time step, the agent will observe a batch of $B=32$ comments. If the agent decides not to publish a comment, a reward of 0.5 is observed regardless of the actual toxicity of the comment. If the agent publishes a non-toxic comment, a reward of 1 is observed. If the agent publishes a toxic comment, a reward of -0.5 is observed. This asymmetric reward function represents a possible example suitable for this real-world scenario. The goal of the agent is to minimize the regret (compared to a clairvoyant optimal publishing policy) over time by learning to make optimal publish/not publish decisions based on the text content.
@@
To investigate the role of uncertainty in LLM bandits, we compare different TS variants with the greedy baseline. Every bandit model is initialized as in Eq. 1, with a pre-trained GPT2 model with 124M parameters (Radford et al., 2019). We use GPT2 (despite larger models being available) because it allows us to perform many experimental runs, making sure our experiments are statistically sound. While this is important in general for LLM research, it is especially vital when it comes to decision-making tasks, where there is inherent stochasticity in the bandit algorithm.
As TS variants, we include dropout, Diagonal Fisher LA (which we will call Diag. LA), LA with full Hessian on the last layer (which we will call Last LA), and Epinet TS. Regarding the epinet architecture, we follow prior work on epinets (Osband et al., 2023a; b) and select an architecture with a multi-layer perceptron $h$ which is multiplied with a dot product with the epistemic index: $\text{epi}_{\eta}(x;z)=h_{\eta}([\text{sg}(\tilde{\pi}_{\theta}(x)),z])^{T}z$. For more details on the architecture, see Appendix B. Every model is trained with regularized MSE loss as in Eq. 3, except for Diag. LA, which updates the prior with the new posterior at each time step, as shown in Eq. 9. All the parameters are updated during training. We train each model at the end of each time step for 50 epochs with the Adam optimizer (Kingma & Ba, 2014), with learning rate set to $3\cdot 10^{-5}.$ For each model, hyperparameters are tuned on 10 random runs (with different seeds than the testing ones) with $T=30$. We did not tune the dropout probability because we wanted to exploit the fact that GPT2 was pre-trained with dropout $p=0.1$ and use the same $p$. We investigate the effectiveness of this choice in the results section. We describe the training phase and the hyperparameter tuning procedure in more detail in Appendix B.

### v2

@@ Abstract
We investigate the role of uncertainty in decision-making problems with natural language as input. For such tasks, using Large Language Models as agents has become the norm. However, none of the recent approaches employ any additional phase for estimating the uncertainty the agent has about the world during the decision-making task. We focus on a fundamental decision-making framework with natural language as input, which is the one of contextual bandits, where the context information consists of text. As a representative of the approaches with no uncertainty estimation, we consider an LLM agent with a greedy policy, which picks the action corresponding to the largest predicted reward. We compare this baseline to LLM agents that make active use of uncertainty estimation by integrating the uncertainty in a Thompson Sampling policy. We employ different techniques for uncertainty estimation, such as Laplace Approximation, Dropout, and Epinets. We empirically show on real-world data that the greedy policy performs worse than the Thompson Sampling policies. These findings suggest that, while overlooked in the LLM literature, uncertainty improves performance on bandit tasks with LLM agents.
@@ 1 Introduction
In parallel, many critical real-world systems are increasingly relying on these models to make decisions (Yang et al., 2023b), where the consequences of a particular action are typically reflected by a reward signal. This usually works by fine-tuning an LLM to serve as a reward model, where we try to teach the network to predict the mean reward for each action. A common way of using this model to make decisions is to use the predictions greedily (Riquelme et al., 2018), taking the action with the greatest estimated reward. However, this approach ignores the fact that reward estimates may be inaccurate, which can lead to never-ending pathological behavior (Russo et al., 2018). On the other hand, the literature on contextual bandits (Wang et al., 2005) provides a principled approach to deal with inaccurate reward estimates in a decision-making task. A popular modelling tool for doing this in the Bayesian framework is to maintain a probabilistic reward model that relies on two separate notions of uncertainty: epistemic and aleatoric.
Epistemic uncertainty is the one that reflects the fact that we have not yet seen enough fine-tuning data to estimate the mean reward well, while aleatoric uncertainty reflects the irreducible noise associated with observing a reward. When such a reward model is coupled with an action-selection algorithm like Thompson Sampling (Thompson, 1933), high-quality estimates of epistemic uncertainty can be leveraged to produce a better balance between exploration and exploitation in the decision-making task, evidenced by lower regret.
In this paper, we want to investigate the importance of employing epistemic uncertainty models in decision-making tasks with natural language and LLMs. In particular, we focus on one of the most fundamental decision-making tasks with natural language as input: contextual bandits. In a contextual bandit problem, the agent must continually take actions based on observed context, which in our case is text, and observes the reward only for the chosen action. This framework is applicable to a myriad of real-world scenarios.
As a concrete example application, let us consider the problem of automated content moderation on online platforms (Gorwa et al., 2020; Vaccaro et al., 2020; Ma & Kou, 2021; Avadhanula et al., 2022). In this scenario, users post comments and the agent decides whether to publish each comment. The user comments serve as the context, in text form. The agent takes an action (e.g. “publish” or “not publish”) based on the context. After acting, the environment provides the reward only for the selected action, which the agent uses to update its policy. For example, if the agent publishes a toxic comment, it receives a large negative reward since other users will report the comment. Instead, if the agent does not publish a comment, it receives a neutral reward regardless of the comment’s content, since the comment is not seen. To learn effectively in this setting, the agent must balance exploration and exploitation.
In this paper, we study the contextual bandit scenario empirically, combining the insights from the LLM and bandit communities. The greedy approach outlined above is our main baseline and was chosen for its simplicity and ease of implementation. We benchmark it against approximate Bayesian models coupled with the Thompson Sampling algorithm for action selection. Since Thompson Sampling requires epistemic uncertainty to be of high quality to work well, we investigate several different approaches of doing so. Specifically, we compare techniques for epistemic uncertainty estimation that can be scaled up to be used with current LLMs. These are the Laplace Approximation (Daxberger et al., 2021), Dropout (Gal & Ghahramani, 2016) and Epinets (Osband et al., 2023a). We show that all our TS policies significantly and consistently outperform the greedy baseline on a real-world bandit benchmark in terms of final regret. A preview of the results is shown in Figure 1. These findings suggest that, while often overlooked in the LLM literature, epistemic uncertainty improves performance on bandit tasks with LLM agents.
@@
  • We provide an extensive empirical study of LLM-based contextual bandits;
@@
  • We empirically show on real-world data that the greedy policy is sub-optimal in terms of final regret, and that epistemic uncertainty is helpful in contextual bandit problems with text data.
@@ 2 Preliminaries
In this section, we describe the problem statement and we introduce the topic of uncertainty in Machine Learning models.
@@ 2.1 Batch Contextual Bandit problem
The contextual bandit problem is a sequential decision-making framework where an agent interacts with an environment over a number of time steps. In particular, we focus on the more general case where, for each time step, we observe a batch of contexts and the agent has to select a batch of actions (Kandasamy et al., 2018).
Let us call the total number of time steps $T$. At each time step $t=1,2,\dots,T$, the agent observes a batch of contexts, which we denote with $x^{1}_{t},\dots,x^{B}_{t}$. Each context $x^{b}_{t}$ contains useful information that should be used to solve the task. In this paper, we focus on decision-making problems driven by text. Hence, in this case, each context $x^{b}_{t}$ consists of text information. The agent then uses the observed data until now, which we call $\mathcal{D}$, to select an action $a^{b}_{t}$ from a set ${\mathcal{A}}$ of $K$ possible actions, conditioned on the given context $x^{b}_{t}$ with $b\in\{1,\dots,B\}$. For each selected action $a^{b}_{t}$, the agent receives a reward $r^{b}_{t}$, that depends on the context $x^{b}_{t}$ and action $a^{b}_{t}$. At the end of the time step, we have observed a set of tuples that we denote as ${\mathcal{D}}_{t}=\{(x^{1}_{t},a^{1}_{t},r^{1}_{t}),\dots,(x^{B}_{t},a^{B}_{t},r^{B}_{t})\}$. This set of tuples is added to the observed dataset $\mathcal{D}\leftarrow\mathcal{D}\cup{\mathcal{D}}_{t}$. The aim of the agent is to learn a policy that minimizes the regret over the $T$ time steps. We define the average regret as the difference between the cumulative reward from the learned policy and the cumulative reward from an optimal policy, all divided by the number of time steps. We define the optimal policy $\pi^{*}$ as the policy that, within the considered policy space $\Pi$, maximizes the expected reward: $\pi^{*}=\arg\max_{\pi\in\Pi}\mathbb{E}_{x}\mathbb{E}_{a\sim\pi(\cdot\mid x)}\left[r(x,a)\right]$. Let us call $r^{*}_{t}$ the expected rewards obtained by the optimal policy at any time step $t$. We can define the average regret as follows:
@@ 3.1 Greedy policy
where $\lambda=\sigma^{2}_{\mathrm{obs}}/\sigma^{2}_{p}$, and we set $\theta^{(0)}=\theta_{p}$. This loss now only assumes that the data points in ${\mathcal{D}}_{t}$ are i.i.d., and we update the prior at each time step with the weights obtained at the previous time step. This will be the loss we will use to optimize the greedy bandit in our experimental analysis in Section 5. Using this loss, we are doing MAP inference many times, applying the Bayes rule anew for each new batch of data. In this way, we perform MAP, which is a well-known established principle in deep learning, while avoiding looping throughout the whole dataset. For a more detailed discussion, we refer to Appendix B. This loss is also present in the literature. For instance, (Daxberger et al., 2021) use this same loss in their continual learning experiment (Daxberger et al. (2021), Appendix C.4.1).
@@ 3.2 Thompson Sampling
Thompson Sampling (TS) (Thompson, 1933; Russo et al., 2018) is a probabilistic algorithm for the contextual bandit problem. The key idea of Thompson Sampling is to maintain an epistemic uncertainty estimate in the form of a posterior distribution over the parameters of the model $P(\theta|{\mathcal{D}})$. Initially, this distribution is set to a prior distribution $P(\theta)$ that represents the agent’s initial uncertainty. On each time step, for each observed context, the agent samples a set of parameters $\hat{\theta}\sim P(\theta|{\mathcal{D}})$ from the posterior distribution, and selects the action with the highest reward according to the sampled model. At the end of the time step, the agent updates the posterior distribution of the parameters with the observed data. A high-level description of this procedure is provided in Algorithm 2.
This approach balances exploration and exploitation by sampling from a posterior that quantifies the epistemic uncertainty. Actions with more uncertain estimates will be explored, while actions that are currently believed to have higher rewards will be exploited. TS can be thought as taking actions according to the (epistemic) probability that they are optimal, which leads to a good balance between exploration and exploitation. The posterior distribution concentrates over time, automatically adjusting the exploration-exploitation trade-off.
Thompson Sampling has been shown to be effective both in theory and in practice (Chapelle & Li, 2011). When the Bayesian updates are exact, there are theoretical guarantees on the performance of TS agents (Agrawal & Goyal, 2017). For instance, Russo & Van Roy (2016) show that Thompson Sampling has a Bayesian regret bound. However, in our case, we are dealing with a complex bandit model that uses a pre-trained LLM. Therefore, an exact Bayesian update is computationally infeasible, and we have to resort to approximations. In Section 4, we will show different techniques to estimate posterior distribution of the parameters of a neural network and how to adapt such techniques to our case of LLM bandits.
@@ 4.1 Dropout
The dropout technique (Srivastava et al., 2014) consists of randomly setting a proportion $p$ of neuron outputs to zero at each forward pass through the network during training. This random “dropping out” of neurons allows the network to sample and train on different (but overlapping) architectures.
In standard supervised learning, dropout is then deactivated during inference, and all dropout neurons are re-scaled to account for the fact that all dropout neurons are active. However, in our case, we still apply dropout during the action selection (Gal & Ghahramani, 2016; Riquelme et al., 2018). Using dropout in this phase, we randomly select (according to the dropout probability) a set of parameters $\hat{\theta}$ to use. This procedure can be seen as sampling parameters from an approximate posterior distribution: $\hat{\theta}\sim P(\theta|{\mathcal{D}})$, and it has connections with variational inference (Gal & Ghahramani, 2016). With such approximate posterior distribution, we can apply Thompson Sampling. This technique is valuable for large models, such as our LLM bandit, because it does not add any overhead to the procedure and does not require additional memory.
@@ 4.2 Laplace Approximation
Notice that the first-order term vanishes because we are expanding around $\theta_{\mathrm{MAP}}$, which is a point of minimum of the loss. After some algebraic manipulations, it can be shown that the posterior distribution can be expressed as [^2] For a more detailed derivation, see Appendix C.:
@@
This approximation has the advantage of being positive semi-definite and (Kunstner et al., 2019) showed that it is accurate when the regression residuals are small. For more details on this approximation, see Appendix C.
@@ 4.3 Last-Layer Laplace Approximation
Alternatively, instead of using Hessian approximations, another way to scale LA for large models is to compute the full Hessian but only for a subset of the parameters of the network, which typically is the last layer of the model. In our case, this means that we will compute the Hessian for the randomly initialized final layer, while the pre-trained LLM layers remain fixed during the sampling phase of TS (every parameter is still fine-tuned during training). This approximation can reduce the exploration due to the fixed weights, but at the same time it allows a full quadratic Hessian computation, leading to a better quality of exploration. Last layer LA is equivalent to Bayesian linear regression (Box & Tiao, 2011) on the features before the last layer within a single batch. Also, with Gaussian likelihood, last-layer LA is exact within a single batch.
@@ 5 Experiments
In this section, we provide an extensive empirical study of LLM-based contextual bandits on real-world data. We show how to adapt the epistemic uncertainty estimation techniques we described to bandits with pre-trained LLMs. In particular, we identify the best variant of the Laplace approximation and dropout to use, in addition to a proposed Epinet architecture that we show works well for TS with LLM bandits. Finally, we empirically show that the greedy policy is sub-optimal (in terms of final regret), shedding light on the importance of epistemic uncertainty for LLM bandits.
@@ 5.1 Experimental Methodology
Tasks
We evaluate the bandit policies on various real-world tasks. These tasks consist of:
  • An open-source dataset, called “measuring hate speech”, which is openly available on HuggingFace Datasets [^4] https://huggingface.co/datasets/ucberkeley-dlab/measuring-hate-speech. This dataset consists of around 136,000 comments. Each of them is associated with a continuous score, called the “hate speech score”, provided for each comment. Comments with a score $>0.5$ are considered “toxic” (around $36\%$ of the comments are labeled as toxic, while the others are not toxic). We will refer to this task as toxic.
  • An open-source dataset, called “IMDb(Maas et al., 2011), which is openly available on HuggingFace Datasets [^5] https://huggingface.co/datasets/imdb. This dataset consists of 50,000 movie reviews. Each of them is associated with a class, which can be “positive” or “negative”, according to the sentiment expressed in the review. The dataset consists of exactly $50\%$ “positive” reviews and $50\%$ “negative” reviews. We will refer to this task as imdb.
  • An open-source dataset, called “Offensive Language Identification(Zampieri et al., 2019), which is openly available on HuggingFace Datasets [^6] https://huggingface.co/datasets/tweet_eval/viewer/offensive. This dataset consists of over 14,000 English tweets. Each of them is classified as “non-offensive” or “offensive”. The dataset consists of $66.9\%$ “non-offensive” tweets and $33.1\%$ “offensive” tweets. We will refer to this task as offensive.
  • An open-source dataset, called “HatEval(Basile et al., 2019), which is openly available on HuggingFace Datasets [^7] https://huggingface.co/datasets/tweet_eval/viewer/hate. This dataset consists of 13,000 English and Spanish tweets, where some of them contain hate speech against immigrants and women. Each of them is classified as “non-hateful” or “hateful”. The dataset consists of $58\%$ “non-hateful” tweets and $42\%$ “hateful” tweets. We will refer to this task as hate.
These tasks are framed as a contextual bandit problem as follows. The context is the text input, and the actions available to the agent are “not publish” or “publish”. For each time step, the agent will observe a batch of $B=32$ comments. The reward function is the following: if the agent decides not to publish a comment, a reward of 0.5 is observed regardless of the actual toxicity of the comment. If the agent publishes a non-toxic comment, a reward of 1 is observed. If the agent publishes a toxic comment, a reward of -0.5 is observed. This asymmetric reward function represents a possible example suitable for this real-world scenario. The goal of the agent is to minimize the regret (compared to a clairvoyant optimal publishing policy) over time by learning to make optimal publish/not publish decisions based on the text content. For the imdb task, we consider “negative” reviews as toxic content; for the offensive task, we consider “offensive” tweets as toxic content; for the hate task, we consider “hateful” tweets as toxic content.
@@
To investigate the role of uncertainty in LLM bandits, we compare different TS variants with the greedy baseline. Every bandit model is initialized as in Eq. 1, with a pre-trained GPT2 model with 124M parameters (Radford et al., 2019). We also include additional experiments with GPT2-XL (1.5B) in order to investigate if our findings generalize to larger models.
As TS variants, we include dropout, Diagonal Fisher LA (which we will call Diag. LA), LA with full Hessian on the last layer (which we will call Last LA), and Epinet TS. Regarding the epinet architecture, we follow prior work on epinets (Osband et al., 2023a; b) and select an architecture with a multi-layer perceptron $h$ which is multiplied with a dot product with the epistemic index: $\text{epi}_{\eta}(x;z)=h_{\eta}([\text{sg}(\tilde{\pi}_{\theta}(x)),z])^{T}z$. For more details on the architecture, see Appendix D. Every model is trained with regularized MSE loss as in Eq. 3, except for Diag. LA, which updates the prior with the new posterior at each time step, as shown in Eq. 9. All the parameters are updated during training. We train each model at the end of each time step for 50 epochs with the Adam optimizer (Kingma & Ba, 2014), with learning rate set to $3\cdot 10^{-5}.$ For each model, hyperparameters are tuned on 10 random runs (with different seeds than the testing ones) with $T=50$ on the toxic task. We did not tune the dropout probability because we wanted to exploit the fact that GPT2 was pre-trained with dropout $p=0.1$ and use the same $p$. We investigate the effectiveness of this choice in the results section. We describe the training phase and the hyperparameter tuning procedure in more detail in Appendix D.